Operations Report Oct 24th

  • Crew 197 Operations Report 24Oct18
    SOL: 3
    Name of person filing report: Jim Ehrhart

Non-nominal systems: solar/batteries
Notes on non-nominal systems: batteries apparently on site, to be
installed Friday
Generator (hours run): 23.5, generator on almost full-time, down 5 min
at approx. 5:30pm to add coolant
Solar— SOC % (Before generator is run at night): N/A
Diesel: 75%
Propane – psi: 45
Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons: Not
checked, ATV not in use

Water (trailer) – gallons: N/A trailer not in use
Water (static) – gallons: estimated 90% full
Trailer to Static Pump used – no or yes: No
Water (loft) – Static to Loft Pump used – yes or no: Yes
Water Meter: morning fill, from 10 to 50, meter 138088.98; evening
fill from 10 to 50, meter 138112.92
Toilet tank emptied: no or yes: Yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Oil Added? No
ATV Fuel Used: Gals No
# Hours the ATVs were Used today: 0
Notes on ATVs: N/A

Deimos rover used: No, out of service
Hours: 134
Beginning charge: N/A
Ending charge: N/A
Currently charging: No
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unk
Beginning charge: Unk
Ending charge: Unk
Currently charging: Unk
Spirit rover used: Yes
Hours: 42.3
Beginning charge: 100
Ending charge: 99
Currently charging: Yes
Opportunity rover used: Yes
Hours: 36.1 hrs
Beginning charge: 100
Ending charge: 100
Currently charging: No
Curiosity rover used: Yes
Hours: 55.3 hrs
Beginning charge: 100
Ending charge: 97
Currently charging: Yes

HabCar used and why, where? No

General notes and comments:

Summary of internet: Internet connectivity and bandwidth fine
Summary of suits and radios: All suits performed nominally. One
charger missing and one wonky – will create two new chargers from
existing spare parts tomorrow with suit engineers. Suits experienced
wear and tear on EVAs, some quilting torn in rover ingress/egress, two
scratches on one helmet
All radios reviewed and settings standardized. 13 units in good shape
and fully charged. 4 units not in good condition and set aside as
Summary of Hab operations: All systems nominal. Fan/heater air filter
replaced. Loft water filter replaced.
Summary of GreenHab operations: Nominal, smoke detector batteries
replaced and tested
Summary of ScienceDome operations: N/A, not in use due to battery issue
Summary of RAMM operations: Nominal, very disorderly, needs cleaning
and organization
Summary of health and safety issues: Health and Safety Officer perform
checks – see report.
Questions, concerns and requests to Mission Support:
– Battery meter needed to check random spares found in RAMM.
– Suggest light in loft

Creative Report Oct 22nd

Crew 197 – Special Report on HAL system – Sol 1 – October 22, 2018

HAL was a system created by Judd Reed, a former Director of Engineering of the MDRS project. It consisted of a laptop with a custom software stack which the early MDRS crews used to monitor their bandwidth, plan out EVAs, check the weather, and write their crew reports in a standard format. It was also intended to automate the crew report process.

Over the years, HAL fell into disuse because it was not being actively maintain, and at some point it stopped working and the laptop was lost. Judd Reed provided me the source code of HAL and I attempted to recruit volunteers to revise and maintain it. Several people suggested starting fresh with a Raspberry Pi-based solution using modern technologies. RPi is a very low cost linux-based computer that can run all modern applications and software stacks.

Starting in 2017, I began working with Josh Baldwin, Matt Hopson and (later) Ryan Watson to create a new HAL solution, including both hardware and software. What we have today is a custom-built Raspberry Pi powered touchscreen that will be wall-mounted and also has a standard monitor, keyboard, and mouse plugged into it.

Our crew and future crews can use the HAL system to write their daily reports and also for general productivity such as web browsing and office documents. The wall-mounted display is intended to show the most important information that the crew needs to be successful including their current bandwidth usage, their daily schedule, current time (Mountain and UTC) down to the second, current weather
conditions, and finally notifications from MarsChat which is a new time-delayed chat application we are also testing out as part of Crew 197.

An early concept image of HAL is visible on this Marspedia web page: https://marspedia.org/MDRS:HAL_Development_Log

Future crew reports will have a photo or two of the HAL system in operation.

– James Burk, Executive Officer, MDRS Crew 197

Sol Summary October 22nd

Crew 197 Daily Summary Report 23.05.2018

MDRS Daily Summary Report for Sol 1

Summary Title
Morning on Mars

Mission Status:

Sol Activity Summary:
5 of us (James B., Shannon N., Marge, Sacha, and Jim E.) arrived from Grand Junction late yesterday, and were greeted warmly by Dr. Rupert as well as NorCal chapter members Max and Robert who delivered the refurbished suits and helmets. We woke up this morning to our first full day on Mars! We are not yet in sim and there was some cleanup and repairs needed to the tunnels between the Hab and the ancillary facilities. Our commander Dr. Susan Ip-Jewell also arrived this afternoon. After dinner, we had an extended training, briefing, and social session with Dr. Rupert.

Look Ahead Plan: It has rained extensively recently, including today, and it’s pretty muddy outside. We are not expecting to do an EVA until Sol 3 (Wednesday). Tomorrow we will get a full briefing of the suits, and plan to enter sim after lunch. Our mission is a short 5 day mission and we hope to survey future drone flights for the MarsVR Program using EVAs, as well as to ensure all of our crew members have great experiences here.

Anomalies in work:
– Tunnels not fully covered with tarps.
– Science Dome is unusable due to battery failures over the
offseason. New batteries will be installed on Friday.
– Mud outside due to rain, hard to do EVAs.

Weather: Nice in the morning. Rainy in the afternoon but has now cleared up.

Crew Physical Status: Everybody is well and in good spirits.

EVA: None planned for tomorrow.

Reports to be filed:

Operations Report by Dr. Rupert

Sol Summary – James Burk (XO)

Hal Report – James Burk (XO)

Crew Photos – James Burk (XO)

Photo of the Day – James Burk (XO)

Support Requested:


Journalist Report October 22nd

Mars Crew 197 – Sol 1 – It’s a Wrap (Already??)

Marge Lipton – Crew 197 Journalist

We arrived at dusk last night and were treated to the soft luscious colors of a southwest sunset. Our crew faced earthly challenges like delayed flights and lost luggage, but persevered, fitting 5 people and too many pieces of luggage to count in a Nissan Pathfinder for the two and a half hour drive to the station. We squeezed everything in, but as newly minted Martians this became a bonding experience. Because it’s clear that besides all the engineering and technology in getting humans to Mars, it’s impossible not to take into consideration how best to get a diverse group of people to get along in such close quarters.

Our first day, today was the start of the MDRS season and involved a lot of housecleaning, from mopping and vacuuming, to setting up new posts to build the simulated pressurized tunnel allowing us to walk between different buildings. After the posts were pounded into the ground, we then WRAPPED the tarps overhead. Rovers had to be moved into position and the MDRS car needed to be taken in for a tune-up.

Crew members discussed what each wanted to accomplish on this mission, and tried to figure out especially, how we were going to remain in sim while also trying to get the best VR coverage of the sites James and Shannon wanted. In practice, this meant, to break sim or not wear a helmet or not on the EVA.

We were interrupted in our WRAPPING of the tunnel by rain, later followed by thunderstorms. So we took the opportunity to make WRAPS of tuna, avocado and tomato.

So as I said, that’s a wrap for today.

Operations Report October 22nd

Operations Report 22OCT2018

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Batteries for power system

Notes on non-nominal systems: Batteries arrive tomorrow and will be installed on Friday!

Generator (hours run): 24

Solar— SOC n/a %

Diesel – 80 %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (axillary tank) – 0 gallons.

Water (static) – 500 gallons

GreenHab water tank: 200 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda used to move Deimos into charging position, the 350 that was not starting needed fuel, because I switched to the reserve tank and it worked fine. This is interesting, because it really did not look empty, it looked about 1/3 fuel. All ATVs, if used, should have a full tank when leaving MDRS. The three 350’s and the 300 have been taken out of sim service. The Honda remains, but only to be used if a rover breaks down and needs tows back to MDRS. Crews will not be trained on the use of ATVs, but one experienced person or chosen person per crew will be trained on how to tow.

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs:

Deimos rover used: Yes, towed into position in the charging bay next to Hab. I will work on it there,


Beginning charge:

Ending charge:

Currently charging: Yes

Sojourner rover used: yes


Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: no.


Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: no


Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used: no


Beginning charge:

Ending charge:

Currently charging: no

HabCar used and why, where? Was used in the am to haul water, then driven to Bicknell for the following repairs: New windshield, new taillight, oil change and safety inspection. We will pick it up at the end of the week and get new tabs for it.

Crew Transport Vehicle: Yes, to drop the HabCar off and to get fence posts for the tunnel

General notes and comments: Work on the tunnel was started today but not finished due to (you guessed it!) rain. Most of the fenceposts were reset and four of the tarps are in place.

Summary of internet: All three accounts nominal. HAL is
operational. More information will be coming about this system.

Summary of suits and radios: Suits have arrived and are being charged for training in the morning.

Summary of Hab operations: All materials have been removed from in front of the Hab except the rocks, which we are referring to as “the in situ resource material”. I will move it over time.

There is a leak in the drain pipe from the summer work done on the plumbing. It is where the old science sink used to drain. For now, we are using the upstairs sink normally, and catching the water in a bucket using the contractor can fix it.

We will be turning on the hot water in the morning and conducting a propane test in the morning.

Summary of GreenHab operations: We began using the big blue water tank in the GreenHab. It is nice to be able to water from inside the building and will provide some temperature buffering capacity.

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of health and safety issues: nothing to report.

Notes, questions, concerns and requests to Mission Support: nothing at this time

Crew 197 arrives at the MDRS in 10 days!

Crew 197 (VR CrowdExplorers)

Overall Goals of Mission

  1. Create a research whitepaper for using VR to train analog simulation participants.
  2. Pilot out MarsVR Training scenarios and script/enact new scenarios.
  3. Update interior/exterior scans for MarsVR future releases.
  4. Capture video & 360 footage for future promotional videos.
  5. Prepare the station for beginning of field season.
    • Installation of new Donor & Memorial plaques. (James)
    • Start up some plants in the GreenHab.
    • Setup of new HAL workstation & displays. (James)
    • Updates to Internet Router firmware & configuration. (James)
    • Other logistical tasks.

Crew will be IN SIM starting the afternoon of Sol 1 (October 22).

Copyright © The Mars Society. All rights reserved. | Main Site